# Statistical Rounding Error Analysis for the Computation of Random Vectors and Matrices

### v2

@@ Abstract
The conventional rounding error analysis provides worst-case bounds with an associated failure probability and ignores the statistical property of the rounding errors. In this paper, we develop a new statistical rounding error analysis for random vectors and matrices computation. By assuming the relative errors are independent random variables, we derive the approximate closed-form expressions for the expectation and variance of the rounding errors in various key computations for vectors and random matrices. Numerical experiments validate the accuracy of our derivations and demonstrate that our analytical expressions are generally at least two orders of magnitude tighter than alternative worst-case bounds, exemplified through the inner products.
@@
Floating-point arithmetic, Random matrix theory, Rounding error analysis, Statistical analysis, Inner products
@@ I Introduction
Rounding error analysis is an essential method for assessing the numerical stability of algorithms, which has been studied extensively in information-theoretic and mathematical literature [1, 2, 3]. Classical rounding error analysis obtains backward error bounds involving the constant $\gamma_{n}=nu/\left(1-nu\right)$ for a dimension $n$ and unit roundoff $u$ [4, Section 3], which offer reasonable backward errors and valuable insights for double-precision arithmetic and moderate dimensions $n$. Nevertheless, the above classical rounding error analysis is very pessimistic about the problem with large dimensions and low-precision arithmetic [5].
@@
Note that the aforementioned works conducted the rounding error analysis from a worst-case bound perspective. The authors in [11] found that the worst-case bounds are very pessimistic for the computation of random vectors and matrices. When the input is a random variable, both the output after rounding and the associated rounding error are also random variables, which is more reasonable to consider the statistical property of rounding errors rather than worst-case bounds. Therefore, if we can conduct the rounding error analysis from a statistical perspective, then we may obtain more accurate and tighter results for the rounding errors of random vectors and matrices computation. Random vectors and matrices computation has numerous applications in the field of wireless communications, signal processing, and machine learning.
@@
In wireless communications, channel matrices are typically treated as random vectors or matrices [12]. Furthermore, precoding and detection processes based on the channel matrix involve random matrix-matrix computations and matrix factorization [13, 14, 15]. In signal processing, random vectors and matrices computation are integral to covariance estimation algorithms [16, 17] and radar waveform design [18, 19]. In addition to wireless communications and signal processing, the computation of random vectors and matrices is also crucial in various machine learning tasks [20, 21].
@@
The above applications motivate the study to give statistical rounding error analysis and derive the statistical property of rounding errors for the computation of random vectors and matrices. There is little research on how to obtain much sharper and more accurate results from a statistical perspective. Note that Constantinides and Dahlqvist [22, 23] derived closed-form expressions for the distribution of the rounding errors under scalar computation when input data are random variables. However, to the best of our knowledge, the expectation and variance of the rounding errors for random vectors and matrices computation are still unknown. Compared to scalar computation, vector and matrix computation involve higher dimensions and different element combinations, leading to the progressive accumulation of rounding errors.
@@
Motivated by the above observations, in this paper, we give a statistical rounding error analysis for random vectors and matrices computation and derive approximate closed-form expressions for the expectation and variance of the rounding errors. Our main contributions are summarized as follows:
@@
  • General rounding error analysis for random vectors and matrices. We analyze rounding errors for the computation of random vectors and matrices with unknown specific distributions from a statistical perspective. Specifically, we derive the expectation and variance of the rounding errors for inner products and provide their approximate closed-form expressions by rigorous proof. Moreover, our analytical expressions can reduce to the probabilistic bounds in [9, 7] based on Bienaymé–Chebyshev inequality. Building on the results for inner products, we extend our analysis to rounding errors in matrix-vector and matrix-matrix products.
@@
Organization: The paper is organized as follows. We first provide the probabilistic floating-point arithmetic model in Sec. II. Then, we apply this model in Sec. III to present the general statistical rounding error for random vectors and matrices. Furthermore, a specific statistical rounding error for Wishart matrices is provided in Sec. IV. In Sec. V, we conduct a series of numerical experiments to validate our analytical expressions thoroughly. We summarize the conclusions in Sec. VI.
Notation: Bold uppercase letters denote matrices and bold lowercase letters denote vectors. For a matrix $\bf A$, ${\bf A}^{T}$, ${\bf A}^{H}$ and ${\bf A}^{-1}$ denote the transpose, the Hermitian transpose and inverse of ${\bf A}$, respectively. ${a}_{ij}$ denotes $(i,j)$th entry of ${\bf A}$. $\mathbb{E}\left(\bf A\right)$ and $\mathbb{V}\left(\bf A\right)$ denotes the expectation and variance of $\bf A$, respectively. $\left|{\bf A}\right|$ represents the matrix of absolute values $(\left|{a}_{ij}\right|)$. Given random variable $a$ and $b$, $\mathbb{C}(a,b)$ are the covariance between $a$ and $b$.
@@ II Floating-point arithmetic model
Note that the bound of the relative error $\delta$ provided by Model 1 is only a worst-case scenario, which will be very pessimistic for numerical analysis [5]. Further, when considering random variables, the statistical property of $\delta$ will become important, since this is the premise for analyzing the statistical characteristics of rounding errors. To end this, the following probabilistic model of relative error is presented.
@@
Let the input signal be independent random variables. In the computation process, the relative errors $\delta$ in Model 1 associated with every pair of operands are independent random variables sampled from a given distribution $\mathcal{DIST}$ and satisfy $\left|\delta\right|\leq u$, where the probability density function (PDF) of $\delta$ is given by [22, 23]
@@
Compared to [8, Model 2], we do not need the input data to be bounded. It is interesting to note from Model 2 that the distribution of the relative error $\delta$ in (3) is approximately deterministic and independent of the input distribution [23]. Furthermore, we can calculate the expectation and variance of $\delta$ using (3) as follows:
@@
Moreover, based on Model 2, we can obtain the probabilistic floating-point arithmetic in the following lemma.
Lemma 1 (Probabilistic floating-point arithmetic [23]).
@@
Based on Model 2 and (4), we observe that the rounding error $\Delta$ is uncorrelated with the input because
@@
$=0=\mathbb{E}\left\{\left(x\,\mathrm{op}\,y\right)\right\}\mathbb{E}\left(\Delta\right).$
Model 2 and corresponding Lemma 1 are not always realistic. For instance, real-world inputs may lack randomness. Additionally, repeated pairs of operands may result in the same $\delta$, i.e., they are dependent. Nevertheless, the question is whether these assumptions effectively model the actual rounding errors encountered in our computations (See similar comments of Hull and Swenson [25] and Kahan [26], as discussed in [7, 8]). We will show that the outcomes obtained under Model 2 closely approximate the actual results through numerical experiments in Sec. V.
Note that many rounding error analyses depend on Model 1, and hence can potentially obtain the distribution (or expectation and variance) of their rounding errors if they can utilize Model 2. Moreover, Lemma 1 only addresses the distribution of rounding errors for scalar computations. When dealing with vector and matrix computations, the complexity and tediousness of calculations make it challenging to derive their rounding error expectation and variance. In the following section, we will provide statistical rounding error analysis for random vectors and matrices based on Model 2.
III General rounding error analysis for random vectors and matrices
In this section, we conduct a general rounding error analysis for random vectors and matrices whose specific distributions are unknown by leveraging Model 2. First, the rounding error analysis for inner products for random vectors is given. Then we extend the rounding error analysis to encompass random matrix-vector and matrix-matrix products.
@@ III General rounding error analysis for random vectors and matrices
Lemma 2 (Expectation and variance of products of random variables [27, Sec. 2]).
@@
Lemma 5 (Expectation and variance of elements in random matrices satisfying Wishart distribution [28, Theroem 3.3.15]).
@@ III-A Rounding Error Analysis for Inner Products
Let ${\bf x},{\bf y}\in\mathbb{R}^{n\times 1}$ be independent random vectors. The entries $x_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{x}$ and variance $\sigma_{x}^{2}$, while the entries $y_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{y}$ and variance $\sigma_{y}^{2}$. If $s={\bf x}^{T}{\bf y}$ is computed in floating-point arithmetic under Model 2, the expectation and variance of the rounding error $\Delta s$ are given by
@@
Let ${\bf x},{\bf y}\in\mathbb{R}^{n\times 1}$ be independent random vectors. The entries $x_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{x}$ and variance $\sigma_{x}^{2}$, while the entries $y_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{y}$ and variance $\sigma_{y}^{2}$. If $s={\bf x}^{T}{\bf y}$ is computed in floating-point arithmetic under Model 2, the computed $\hat{s}$ stratifies
@@
Corollary 1 shows that we can obtain a probabilistic $\mathcal{O}(\sqrt{n}u)$ bound in (10) (similar to the bound in [7]) from the derived result in Theorem 1 for random vectors. Moreover, when random vectors have a mean of zero, the backward error bound in (11) is of order $\mathcal{O}(u)$ and does not grow $n$ (similar to the analysis in [8]). Therefore, the derived results in Theorem 1 are general and can reduce to some classic bounds in [7, Eq. (3.6)] and [8, Eq. (3.3)].
@@ IV Specific rounding error analysis for Wishart Matrices
One traditional approach for solving the LS problem is the normal equation method [33, Alg. 5.3.1]. Specifically, we can use the following procedure:
@@
where ${\bf A}$ is a Wishart matrix satisfying the Wishart distribution $\mathcal{W}_{n}\left(m,{\bf I}_{n}\right)$. It is observed that the standard algorithms such as matrix-matrix products, matrix-vector products, Cholesky factorization, LU factorization, and the solution of triangular systems are used in (18). Since the rounding error analysis for matrix-matrix and matrix-vector multiplication has been presented in Sec. III, we further derive the rounding error analysis for the solution of triangular systems and matrix factorization under the condition of Wishart matrices.
@@ IV-A Rounding Error Analysis for The Solution of Triangular Systems
In the subsection, we provide the rounding error analysis for the solution of triangular systems. Without loss of generality, a lower triangular matrix is considered. Specifically, let the lower triangular matrix ${\bf T}=\left(t_{ij}\right)$ be obtained through the Cholesky factorization of Wishart matrix ${\bf A}\sim\mathcal{W}_{n}\left(m,{\bf I}_{n}\right)$ in (18). Furthermore, we have $t_{ij},1\leq j\leq i\leq n$ are independently distributed, $t_{ii}^{2}\sim\chi_{m-i+1}^{2},1\leq i\leq n$ and $t_{ij}\sim\mathcal{N}\left(0,1\right),1\leq j<i\leq n$ [28, Theorem 3.3.4].
Then, given a triangular system ${\bf Tx}={\bf b}$, the solution $\bf x$ can be computed as follows: [34, Algorithm 3.1.3]
@@ IV-B Rounding Error Analysis for LU Factorization
In the subsection, we present the rounding error analysis for LU factorization of Wishart matrix ${\bf A}\sim\mathcal{W}_{n}\left(m,{\bf I}_{n}\right)$. Specifically, the LU factors of ${\bf A}$ are given by the Doolittle form of Gaussian elimination in the following recurrences [4, Algorithm 9.2]:
@@
    $f_{u_{ij}}\left(z\right)=\begin{cases}\frac{\Gamma\left(\frac{\nu-1}{2}\right)}{2\sqrt{\pi}\Gamma\left(\frac{\nu}{2}\right)},&0\\ \frac{1}{\sqrt{2\pi}2^{\frac{\nu}{2}-1}\Gamma\left(\frac{\nu}{2}\right)}\left(\left|z\right|\right)^{\frac{\nu-1}{2}}K_{\frac{\nu-1}{2}}\left(\left|z\right|\right),&else\\ \end{cases},$
@@
For the Cholesky factorization ${\bf A}={\bf R}^{T}{\bf R}$, where ${\bf R}=\left(r_{ij}\right)$ is a upper triangular matrix with $r_{ii}>0$, we have $r_{ij},1\leq i\leq j\leq n$ are independently distributed, $r_{ii}^{2}\sim\chi_{m-i+1}^{2},1\leq i\leq n$ and $r_{ij}\sim\mathcal{N}\left(0,1\right),1\leq i<j\leq n$ [28, Theorem 3.3.4]. Note that the relationship between LU factorization and Cholesky factorization can be established through the $LDL^{T}$ factorization, i.e.,
@@
where $(a)$ follows [35, Eq. (3.478.4)], and $K_{n}\left(y\right)$ is the modified Bessel function of the second kind. Moreover, note that $K_{n}\left(y\right)\longrightarrow\infty$ when $y\longrightarrow 0$. We can use the asymptotic form for small arguments of the Bessel function [36] and have
@@
  • Denote $q_{i}=l_{ki}u_{ij},1\leq i\leq k-1,k\leq j\leq n,1\leq k\leq n$. If $j=k$, we have
    $\mathbb{E}\left(q_{i}\right)$ $=1,~{}\mathbb{V}\left(q_{i}\right)=2,~{}\mathbb{C}\left(q_{s},q_{t}\right)=0,~{}s\neq t,$
    $\mathbb{C}\left(a_{kk},q_{i}\right)$ $=\mathbb{E}\left(a_{kk}q_{i}\right)=2.$
@@
    $\mathbb{E}\left(q_{i}\right)$ $=0,~{}\mathbb{V}\left(q_{i}\right)=1,~{}\mathbb{C}\left(q_{s},q_{t}\right)=0,~{}s\neq t,$
    $\mathbb{C}\left(a_{kj},q_{i}\right)$ $=\mathbb{E}\left(a_{kj}q_{i}\right)=1.$
  • Denote $p_{j}=\frac{l_{ij}u_{jk}}{u_{kk}},k+1\leq i\leq n,1\leq j\leq k-1,1\leq k\leq n$. We have
    $\mathbb{E}\left(p_{j}\right)$ $=0,~{}\mathbb{V}\left(p_{j}\right)=\frac{1}{\left(m-k-1\right)\left(m-k-3\right)},$
    $\mathbb{C}\left(p_{s},p_{t}\right)$ $=0.$
  • Denote $o_{i}=\frac{a_{ik}}{u_{kk}},k+1\leq i\leq n,1\leq j\leq k-1,1\leq k\leq n$. We have
    $\mathbb{E}\left(o_{i}\right)$ $=0,~{}\mathbb{V}\left(o_{i}\right)=\frac{m-4}{\left(m-k-1\right)\left(m-k-3\right)},$
    $\mathbb{C}\left(o_{i},p_{j}\right)$ $=\mathbb{E}\left(o_{i}p_{j}\right)=\frac{1}{\left(m-k-1\right)\left(m-k-3\right)}.$
@@
From the proof of Lemma 7, we derive $l_{ij}=\frac{r_{ji}}{r_{jj}}$ and $u_{ij}=r_{ii}r_{ij}$. Subsequently, we can obtain $q_{i}=r_{ik}r_{ij}$, $p_{j}=\frac{r_{ji}r_{jk}}{r_{kk}^{2}}$, and $o_{i}=\frac{r_{ki}+\sum_{j=1}^{k-1}{r_{ji}r_{jk}}}{r_{kk}}$. Note that the distribution of $r_{ij}$ is already known, thus concluding the proof. ∎

### v3

# Statistical Rounding Error Analysis for Random Matrix Computations

@@ Abstract
The conventional rounding error analysis provides worst-case bounds with an associated failure probability and ignores the statistical property of the rounding errors. In this paper, we develop a new statistical rounding error analysis for random matrix computations. Such computations occur, e.g., during the precoding or detection process for massive multiple-input multiple-output (MIMO) systems. By assuming the relative errors are independent random variables, we derive the approximate closed-form expressions for the expectation and variance of the rounding errors in various key computations for random matrices. Numerical experiments validate the accuracy of our derivations and demonstrate that our analytical expressions are generally at least two orders of magnitude tighter than alternative worst-case bounds, exemplified through the inner products.
@@
Floating-point arithmetic, Matrix computations, Random matrices, Rounding error analysis
@@ I Introduction
Random matrix computations problems have been studied extensively in signal processing literature [1, 2, 3]. One of the central questions in the study of random matrix computations is to analyze the rounding error [4, 5, 6, 7]. Classical rounding error analysis obtains backward error bounds involving the constant $\gamma_{n}=nu/\left(1-nu\right)$ for a dimension $n$ and unit roundoff $u$ [8, Chapter 3], which offer reasonable backward errors and valuable insights for double-precision arithmetic and moderate dimensions $n$. Nevertheless, the above classical rounding error analysis is very pessimistic about the problem with large dimensions and low-precision arithmetic [9].
@@
Note that the aforementioned works conducted the rounding error analysis from a worst-case bound perspective. The authors in [15] applied the above worst-case bounds to the precoding and detection processes and found them to be overly pessimistic. This is because the precoding and detection processes involve random matrix computations. In this case, worst-case scenarios are often rare from a statistical perspective. In other words, when the input is a random variable, both the output after rounding and the associated rounding error are also random variables, which is more reasonable to consider the statistical property of rounding errors rather than worst-case bounds. Therefore, if we can conduct the rounding error analysis from a statistical perspective, then we may obtain more accurate and tighter results for the rounding errors of random matrix computations. Random matrix computations have numerous applications in the field of wireless communications, signal processing, and machine learning.
@@
In wireless communications, channel matrices are typically treated as random vectors or matrices [16]. Furthermore, precoding and detection processes based on the channel matrix involve random matrix-matrix computations and matrix factorization [17, 18, 19]. In signal processing, random matrix computations are integral to covariance estimation algorithms [20, 21] and radar waveform design [22, 23]. In addition to wireless communications and signal processing, the computation of random matrices is also crucial in various machine learning tasks [24, 25].
@@
The above applications motivate the study to give statistical rounding error analysis and derive the statistical property of rounding errors for the computation of random matrices. There is little research on how to obtain much sharper and more accurate results from a statistical perspective. Note that Constantinides et al. [26] and Dahlqvist et al. [27] derived closed-form expressions for the distribution of the rounding errors under scalar computation when input data are random variables. However, to the best of our knowledge, the expectation and variance of the rounding errors for random matrix computations are still unknown. Compared to scalar computation, vector and matrix computation involve higher dimensions and different element combinations, leading to the progressive accumulation of rounding errors.
@@
Motivated by the above observations, in this paper, we give a statistical rounding error analysis for random matrix computations and derive approximate closed-form expressions for the expectation and variance of the rounding errors. Our main contributions are summarized as follows:
@@
  • General rounding error analysis for random matrices. We analyze rounding errors for the computation of random matrices with unknown specific distributions from a statistical perspective. Specifically, we derive the expectation and variance of the rounding errors for inner products and provide their approximate closed-form expressions by rigorous proof. Moreover, our analytical expressions can reduce to the probabilistic bounds in [13, 11] based on Bienaymé–Chebyshev inequality. Building on the results for inner products, we extend our analysis to rounding errors in matrix-vector and matrix-matrix products.
@@
Organization: The paper is organized as follows. We first provide the probabilistic floating-point arithmetic model in Section II. Then, we apply this model in Section III to present the general statistical rounding error for random matrices. Furthermore, a specific statistical rounding error for Wishart matrices is provided in Section IV. In Section V, we conduct a series of numerical experiments to validate our analytical expressions thoroughly. We summarize the conclusions in Section VI.
Notation: Bold uppercase letters denote matrices and bold lowercase letters denote vectors. For a matrix $\bf A$, ${\bf A}^{T}$, ${\bf A}^{H}$ and ${\bf A}^{-1}$ denote the transpose, the Hermitian transpose and inverse of ${\bf A}$, respectively. ${a}_{ij}$ denotes $(i,j)$th entry of ${\bf A}$. $\mathbb{E}\left(\cdot\right)$ and $\mathbb{V}\left(\cdot\right)$ denotes the expectation and variance, respectively. $\left|{\bf A}\right|$ represents the matrix of absolute values $(\left|{a}_{ij}\right|)$. Given random variable $a$ and $b$, $\mathbb{C}(a,b)$ are the covariance between $a$ and $b$.
@@ II Floating-point arithmetic model
Model 1 illustrates that finite-precision arithmetic operations can introduce a relative error $\delta$ for some $\delta\in[-u,u]$. In other words, Model 1 is non-deterministic, and the specific value of $\delta$ in (2) is unknown. This means that numerical analysis based on Model 1 must consider all possible values $\delta$, i.e. it is based on Model 1 is fundamentally a worst-case analysis. Notably, when the inputs are random variables (i.e., probabilistic), both the output after rounding and the associated relative error $\delta$ become random variables from a statistical perspective. In such cases, Model 1 may be overly conservative, as worst-case scenarios are often rare and may not occur in practice. To this end, the following probabilistic model of relative error and probabilistic floating-point arithmetic is provided.
@@
Let the input signal be independent random variables. In the computation process, the relative errors $\delta$ in Model 1 associated with every pair of operands are assumed to be independent random variables sampled from a given distribution $\mathcal{DIST}$ and satisfy $\left|\delta\right|\leq u$, where the probability density function (PDF) of $\delta$ is given by [26, 27]
@@
Compared to [12, Model 2], we do not need the input data to be bounded. It is interesting to note from Model 2 that the distribution of the relative error $\delta$ in (3) is approximately deterministic and is assumed to be independent of the input distribution. Furthermore, we can calculate the expectation and variance of $\delta$ using (3) as follows:
@@
Moreover, based on Model 2, we can obtain the probabilistic floating-point arithmetic in the following definition.
Definition 1 (Probabilistic floating-point arithmetic [27]).
@@
Based on Model 2 and (4), we observe that the rounding error $\Delta$ is approximately uncorrelated with the input because
@@
$\approx 0\approx\mathbb{E}\left\{\left(x\,\mathrm{op}\,y\right)\right\}\mathbb{E}\left(\Delta\right).$
Model 2 and corresponding Definition 1 are not always realistic. For instance, real-world inputs may lack randomness. Additionally, repeated pairs of operands may result in the same $\delta$, i.e., they are dependent. Nevertheless, the question is whether these assumptions effectively model the actual rounding errors encountered in our computations (See similar comments of Hull and Swenson [29] and Kahan [30], as discussed in [11, 12]). We will show that the outcomes obtained under Model 2 closely approximate the actual results through numerical experiments in Section V. Furthermore, numerical experiments show that Model 2 is valid for independent random variables.
Note that many rounding error analyses depend on Model 1, and hence can potentially obtain the distribution (or expectation and variance) of their rounding errors if they can utilize Model 2. Moreover, Definition 1 only addresses the distribution of rounding errors for scalar computations. When dealing with vector and matrix computations, the complexity and tediousness of calculations make it challenging to derive their rounding error expectation and variance.
Overall, our goal in the following sections is to derive closed-form expressions of expectation and variance for the computation of random matrices based on Model 2 and to validate the accuracy of these expressions.
III General rounding error analysis for random matrices
In this section, we conduct a general rounding error analysis for random matrices whose specific distributions are unknown by leveraging Model 2. First, the rounding error analysis for inner products for random vectors is given. Then we extend the rounding error analysis to encompass random matrix-vector and matrix-matrix products.
@@ III General rounding error analysis for random matrices
Lemma 1 (Expectation and variance of products of random variables [31, Section 2]).
@@
Lemma 4 (Expectation and variance of elements in random matrices satisfying Wishart distribution [32, Theroem 3.3.15]).
@@ III-A Rounding Error Analysis for Inner Products
Let ${\bf x},{\bf y}\in\mathbb{R}^{n\times 1}$ be independent random vectors. The entries $x_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{x}$ and variance $\sigma_{x}^{2}$, while the entries $y_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{y}$ and variance $\sigma_{y}^{2}$. If $s={\bf x}^{T}{\bf y}$ is computed in floating-point arithmetic under Model 2, the expectation and variance of the rounding error $\Delta s$ are given in (6) and (7) at the bottom of the page.
@@
Let ${\bf x},{\bf y}\in\mathbb{R}^{n\times 1}$ be independent random vectors. The entries $x_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{x}$ and variance $\sigma_{x}^{2}$, while the entries $y_{i}$ for $i=1$ to $n$ are sampled from a distribution with mean $\mu_{y}$ and variance $\sigma_{y}^{2}$. If $s={\bf x}^{T}{\bf y}$ is computed in floating-point arithmetic under Model 2, the computed $\hat{s}$ satisfies
@@
Corollary 1 shows that we can obtain a probabilistic $\mathcal{O}(\sqrt{n}u)$ bound in (10) (similar to the bound in [11]) from the derived result in Theorem 1 for random vectors. Moreover, when random vectors have a mean of zero, the backward error bound in (11) is of order $\mathcal{O}(u)$ and does not grow $n$ (similar to the analysis in [12]). Therefore, the derived results in Theorem 1 are general and can reduce to some classic bounds in [11, Eq. (3.6)] and [12, Eq. (3.3)].
@@ IV Specific rounding error analysis for Wishart Matrices
One traditional approach for solving the LS problem is the normal equation method [37, Alg. 5.3.1]. Specifically, we can use the following procedure:
@@
where ${\bf A}$ is a Wishart matrix satisfying the Wishart distribution $\mathcal{W}_{n}\left(m,{\bf I}_{n}\right)$. It is observed that the standard algorithms such as matrix-matrix products, matrix-vector products, Cholesky factorization, LU factorization, and the solution of triangular systems are used in (18). Since the rounding error analysis for matrix-matrix and matrix-vector multiplication has been presented in Section III, we further derive the rounding error analysis for the solution of triangular systems and matrix factorization under the condition of Wishart matrices.
@@ IV-A Rounding Error Analysis for The Solution of Triangular Systems
In the subsection, we provide the rounding error analysis for the solution of triangular systems. Without loss of generality, a lower triangular matrix is considered. Specifically, let the lower triangular matrix ${\bf T}=\left(t_{ij}\right)$ be obtained through the Cholesky factorization of Wishart matrix ${\bf A}\sim\mathcal{W}_{n}\left(m,{\bf I}_{n}\right)$ in (18). Furthermore, we have $t_{ij},1\leq j\leq i\leq n$ are independently distributed, $t_{ii}^{2}\sim\chi_{m-i+1}^{2},1\leq i\leq n$ and $t_{ij}\sim\mathcal{N}\left(0,1\right),1\leq j<i\leq n$ [32, Theorem 3.3.4].
Then, given a triangular system ${\bf Tx}={\bf b}$, the solution $\bf x$ can be computed as follows: [38, Algorithm 3.1.3]
@@ IV-B Rounding Error Analysis for LU Factorization
In the subsection, we present the rounding error analysis for LU factorization of Wishart matrix ${\bf A}\sim\mathcal{W}_{n}\left(m,{\bf I}_{n}\right)$. Specifically, the LU factors of ${\bf A}$ are given by the Doolittle form of Gaussian elimination in the following recurrences [8, Algorithm 9.2]:
@@
    $f_{u_{ij}}\left(z\right)=\begin{cases}\frac{\Gamma\left(\frac{\nu-1}{2}\right)}{2\sqrt{\pi}\Gamma\left(\frac{\nu}{2}\right)},&z=0\\ \frac{1}{\sqrt{2\pi}2^{\frac{\nu}{2}-1}\Gamma\left(\frac{\nu}{2}\right)}\left(\left|z\right|\right)^{\frac{\nu-1}{2}}K_{\frac{\nu-1}{2}}\left(\left|z\right|\right),&else\\ \end{cases},$
@@
For the Cholesky factorization ${\bf A}={\bf R}^{T}{\bf R}$, where ${\bf R}=\left(r_{ij}\right)$ is a upper triangular matrix with $r_{ii}>0$, we have $r_{ij},1\leq i\leq j\leq n$ are independently distributed, $r_{ii}^{2}\sim\chi_{m-i+1}^{2},1\leq i\leq n$ and $r_{ij}\sim\mathcal{N}\left(0,1\right),1\leq i<j\leq n$ [32, Theorem 3.3.4]. Note that the relationship between LU factorization and Cholesky factorization can be established through the $LDL^{T}$ factorization, i.e.,
@@
where $(a)$ follows [39, Eq. (3.478.4)], and $K_{n}\left(y\right)$ is the modified Bessel function of the second kind. Moreover, note that $K_{n}\left(y\right)\longrightarrow\infty$ when $y\longrightarrow 0$. We can use the asymptotic form for small arguments of the Bessel function [40] and have
@@
$\small\mathbb{E}\left(\Delta u_{kk}\right)$ $=0,$ (25)
$\mathbb{V}\left(\Delta u_{kk}\right)$ $\approx\left(m^{2}-4\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]-3\left(k-1\right)+3\sum_{i=1}^{k-1}{\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}}$
$\mathbb{E}\left(\Delta u_{kj}\right)$ $=0,$ (27)
$\mathbb{V}\left(\Delta u_{kj}\right)$ $\approx\left(m-2\right)\left[\left(1+\sigma^{2}\right)^{k-1}-1\right]+\sum_{i=1}^{k-1}{\left(1+\sigma_{\epsilon_{i}}^{2}\right)\left(1+\sigma_{\eta_{i}}^{2}\right)\left(1+\sigma^{2}\right)^{k-i+1}}-2\frac{\left(1+\sigma^{2}\right)\left[\left(1+\sigma^{2}\right)^{k-2}-1\right]}{\sigma^{2}}+k-3,$ (28)
$\mathbb{E}\left(\Delta l_{ik}\right)$ $=0,$ (29)
$\mathbb{V}\left(\Delta l_{ik}\right)$ $\approx\frac{(m-6)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}+\frac{\left(1+\sigma_{\eta_{k}}^{2}\right)\sum_{j=1}^{k-1}{\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}}-k+1}{\left(m-k-1\right)\left(m-k-3\right)}$
@@
  • Denote $q_{kij}=l_{ki}u_{ij},1\leq i\leq k-1,k\leq j\leq n,1\leq k\leq n$. If $j=k$, we have
    $\mathbb{E}\left(q_{kij}\right)$ $=1,~{}\mathbb{V}\left(q_{kij}\right)=2,~{}\mathbb{C}\left(q_{ksj},q_{ktj}\right)=0,~{}s\neq t,$
    $\mathbb{C}\left(a_{kk},q_{kij}\right)$ $=\mathbb{E}\left(a_{kk}q_{kij}\right)=2.$
@@
    $\mathbb{E}\left(q_{kij}\right)$ $=0,~{}\mathbb{V}\left(q_{kij}\right)=1,~{}\mathbb{C}\left(q_{ksj},q_{ktj}\right)=0,~{}s\neq t,$
    $\mathbb{C}\left(a_{kj},q_{kij}\right)$ $=\mathbb{E}\left(a_{kj}q_{kij}\right)=1.$
  • Denote $p_{ijk}=\frac{l_{ij}u_{jk}}{u_{kk}},k+1\leq i\leq n,1\leq j\leq k-1,1\leq k\leq n$. We have
    $\mathbb{E}\left(p_{ijk}\right)$ $=0,~{}\mathbb{V}\left(p_{ijk}\right)=\frac{1}{\left(m-k-1\right)\left(m-k-3\right)},$
    $\mathbb{C}\left(p_{isk},p_{itk}\right)$ $=0.$
  • Denote $o_{ik}=\frac{a_{ik}}{u_{kk}},k+1\leq i\leq n,1\leq j\leq k-1,1\leq k\leq n$. We have
    $\mathbb{E}\left(o_{ik}\right)$ $=0,~{}\mathbb{V}\left(o_{ik}\right)=\frac{m-4}{\left(m-k-1\right)\left(m-k-3\right)},$
    $\mathbb{C}\left(o_{ik},p_{ijk}\right)$ $=\mathbb{E}\left(o_{ik}p_{ijk}\right)=\frac{1}{\left(m-k-1\right)\left(m-k-3\right)}.$
@@
From the proof of Lemma 6, we derive $l_{ij}=\frac{r_{ji}}{r_{jj}}$ and $u_{ij}=r_{ii}r_{ij}$. Subsequently, we can obtain $q_{kij}=r_{ik}r_{ij}$, $p_{ijk}=\frac{r_{ji}r_{jk}}{r_{kk}^{2}}$, and $o_{ik}=\frac{r_{ki}+\sum_{j=1}^{k-1}{r_{ji}r_{jk}}}{r_{kk}}$. Note that the distribution of $r_{ij}$ is already known, thus concluding the proof. ∎